OpenStudy (anonymous):

which rectangular equation is equivalent to the polar equation r csc theta=8 A. y=8 B. x^2+(y-4)^2=16 C. x=8y^2 D. x=8

OpenStudy (nurali):

No idea i think the option is B

OpenStudy (asnaseer):

do you know what csc(theta) is?

OpenStudy (anonymous):

1/sin theta?

OpenStudy (asnaseer):

good. so your equation is:\[\frac{r}{\sin(\theta)}=8\]so, first multiply both sides by \(\sin(\theta)\) - what do you end up with?

OpenStudy (anonymous):

r=8sintheta

OpenStudy (asnaseer):

great! now the next "trick" to notice is that if you now multiply both sides by "r" you should get some familiar terms - what do you get?

OpenStudy (anonymous):

8sintheta(r) ?

OpenStudy (asnaseer):

both sides

OpenStudy (anonymous):

8sinthetar=8sintheta r ? like that?

OpenStudy (asnaseer):

no, you currently have: r = 8sin(theta) now multiply both sides of this equation by r - what will you end up with?

OpenStudy (anonymous):

thats why doesnt the r cancel out and moves to the other side of the euqation. 8sin(theta)r

OpenStudy (asnaseer):

we are "multiplying" not dividing - so the "r" should not cancel out

OpenStudy (anonymous):

so its r^2?

OpenStudy (asnaseer):

yes - that is what you get on the left-hand-side of the equals sign, and what about the right-hand-side?

OpenStudy (anonymous):

r^2=8 sin(theta)r

OpenStudy (asnaseer):

great! this is usually written as:\[r^2=8r\sin(\theta)\]now, can you spot what to do next?

OpenStudy (anonymous):

square root both sides to cancel out the squared?